OpenStudy (anonymous):

i do not know how to do this. one base of a trapezoid is twice as long as the other. the height is the average of the 2 bases. the area is 324 cm squared. find the height and lengths of the bases.

OpenStudy (anonymous):

A=(h(b1+b2))/2 b1=x, b2=2s, h=b1+b2=3x 324=(3x(2x+x))/2 324=9x/2 (324 x 2)/9=x x=72,subtitution find base, height base 1=72 , base 2= 144, height=216 check: (216 x (72+144))/2= (216 x216)/2 A=324

OpenStudy (anonymous):

A=(h(b1+b2))/2 b1=x, b2=2s, h=b1+b2=3x 324=(3x(2x+x))/2 324=9x/2 (324 x 2)/9=x x=72,subtitution find base, height base 1=72 , base 2= 144, height=216 check: (216 x (72+144))/2= (216 x216)/2 A=324

OpenStudy (anonymous):

A=(h(b1+b2))/2 b1=x, b2=2s, h=b1+b2=3x 324=(3x(2x+x))/2 324=9x/2 (324 x 2)/9=x x=72,subtitution find base, height base 1=72 , base 2= 144, height=216 check: (216 x (72+144))/2= (216 x216)/2 A=324

OpenStudy (anonymous):

A=(h(b1+b2))/2 b1=x, b2=2s, h=b1+b2=3x 324=(3x(2x+x))/2 324=9x/2 (324 x 2)/9=x x=72,subtitution find base, height base 1=72 , base 2= 144, height=216 check: (216 x (72+144))/2= (216 x216)/2 A=324